# A Measurement Study of Model Context Protocol

### v1

@@ 6.3. Measurement of Functionality and Implementation for Exposure Analysis
To comprehensively understand the MCP ecosystem from a security perspective, we measure servers along two functional dimensions: categories of exposed services, and implementation languages. Each of these dimensions provides a distinct lens on potential risks. First, server categories matter because certain functionalities such as authentication, personal data connectors, or proprietary model access, are inherently security-sensitive and expand the attack surface. Second, the implementation language is closely tied to security posture, since different language vary in their unsafe defaults (e.g., memory safety issues in C/C++ vs. stronger isolation in Rust).
@@
Answer to RQ2 (Security and Privacy Posture). Our measurements show that the MCP ecosystem is neither uniformly secure nor privacy-preserving, but instead exhibits a mix of encouraging practices and systemic risks. On the positive side, a large fraction of servers remain actively maintained (40.9% updated within the last 90 days), and many projects incorporate safeguards such as schema validation frameworks (e.g., pydantic, zod) that reduce the likelihood of injection or deserialization flaws. However, multiple categories of servers inherently handle sensitive user data, where misconfigurations or weak access controls could directly expose personal records, documents, or API tokens. The language distribution amplifies this risk: more than 93% of MCP servers are implemented in JavaScript or Python, creating a supply-chain monoculture where vulnerabilities in popular npm or PyPI packages could cascade across thousands of deployments.
@@ 7. Measurement of MCP Clients
In this section, we conduct a systematic analysis of MCP clients, which function as the crucial interface between end-users, MCP servers, and the underlying LLMs that power model-driven tasks. By design, clients mediate the flow of requests and responses, orchestrating communication and shaping the overall user experience. Our measurement focuses on two complementary aspects. First, we examine communication protocols, which uncover whether client–server interactions are converging on a dominant mode or remain fragmented across competing designs. Protocol choice is not only a technical matter but also a signal of standardization, interoperability, and long-term ecosystem trajectory. Second, we analyze cross-server usage, capturing whether clients are designed for single-purpose connections or instead support multi-server integration, which enables richer workflows, redundancy, and interoperability across heterogeneous backends. Together, these two perspectives highlight the balance between simplicity and extensibility in client design, and how this balance contributes to the ecosystem’s evolutionary path. To support this analysis, MCPCrawler collected a total of 341 valid MCP clients across four major markets. As part of a preprocessing step, we applied strict filtering criteria to remove 22 low-value entries, such as inactive forks, placeholder projects, or duplicates that added noise but no substantive functionality.
@@ 7.1. Measurement of Connection Protocols for Standardization Analysis
Figure 12. Interaction modes of MCP clients.
To characterize client connection patterns, we collected metadata from 341 valid MCP clients, extracting their declared interaction protocols. We then normalized protocol labels (e.g., SSE, stdio, others) and computed their relative frequencies. The results are visualized in Figure 12 as a pie chart, where each slice corresponds to the proportion of clients adopting a given protocol. This representation provides an intuitive view of the dominant interaction modes across the ecosystem. The measurement shows that SSE dominates (56.98%, 194 clients), followed by stdio (38.12%, 130 clients), while other protocols remain marginal (4.99%, 17 clients). This skewed distribution suggests that the ecosystem is already converging toward SSE as the preferred interaction model, reinforcing its role as a candidate standard for future interoperability. At the same time, the persistence of stdio, with more than one-third of clients, indicates that alternative design philosophies remain relevant, especially in lightweight or local contexts. The coexistence of these modes highlights a transitional phase in the ecosystem’s evolution: while SSE is becoming the mainstream, diversity in connection patterns still provides space for experimentation and specialized use cases.
@@ 7.2. Measurement of Client Connection Modes for Usage Dynamics
Answer to RQ3 (Client Connection Patterns in MCP Evolution). Our measurements of client protocols and connection modes together show how MCP client design is shaping the trajectory of ecosystem adoption and standardization. On the protocol side, the majority of clients (56.9%) already rely on SSE, signaling a convergence toward this mode as the de facto standard for interoperability. At the same time, a sizable minority (38.1%) still use stdio, underscoring that lightweight, local workflows remain relevant and that the ecosystem is still in a transitional phase where multiple design choices coexist. On the connection side, most clients (80.9%) support only a single server link, favoring simplicity and one-to-one usage, while nearly one-fifth (19.1%) enable multi-server integration, pointing to emerging needs for richer workflows, redundancy, and cross-service aggregation. These results indicate that while MCP clients are converging on SSE as the dominant interaction protocol, diversity persists both in protocol choice and in connection multiplicity. This combination reflects an ecosystem that is gradually moving toward standardization but still leaves room for flexibility, experimentation, and integration-oriented evolution.
@@ 8. Discussion
Ethical Considerations. Our study was carried out with explicit attention to research ethics. All data we analyzed were drawn from publicly accessible sources, including open repositories, market-facing entries, and metadata exposed through official project listings. At no point did we attempt to bypass authentication mechanisms, crawl private datasets, or interact with servers in ways that could degrade their availability or stability. Importantly, downloading metadata and dependency information from GitHub or other MCP markets does not diminish the commercial or functional value of those platforms; rather, it mirrors the type of visibility already available to any developer or user browsing these resources. We did not attempt to probe or exploit server endpoints, ensuring that our methodology posed no threat to operational security. Beyond technical measures, we recognize broader ethical questions around ecosystem measurement, particularly the potential for exposing weaknesses that malicious actors could exploit. To mitigate this, our presentation emphasizes aggregate trends and structural observations, rather than detailed exploit paths.
@@ 9. Related Work
The MCP and the broader ecosystem of LLM plugin frameworks are emerging research frontiers that blend protocol design, marketplace dynamics, and security considerations. Because MCP itself was only recently introduced, existing literature remains sparse and fragmented. To contextualize our measurement study, we review three closely related lines of work: (i) early analyses of the MCP ecosystem, and (ii) open-source agent frameworks and tool libraries that share conceptual similarities with MCP but differ in deployment and measurement scope.
MCP Ecosystem Analyses. The Model Context Protocol (MCP) is a recent open JSON-RPC standard for connecting large language models (LLMs) to external tools and data sources (anthropic2024mcp, ). To date, only a handful of academic studies have investigated MCP. Li et al. (li2025mcpmeasurement, ) conduct one of the first empirical analyses of real-world MCP-based plugins, revealing that network and system APIs dominate MCP servers and that low-download plugins frequently embed high-risk calls. Their work highlights the potential of MCP as a unifying protocol for LLM–tool integration but does not report deployment statistics or adoption rates. Similarly, Anthropic’s official technical documentation (anthropic2024mcp, ) defines the specification and server lifecycle but provides no usage analytics. Parallel to MCP, several works have analyzed OpenAI’s plugin-based ecosystems. Iqbal et al. (iqbal2024llmplatformsecurityapplying, ) propose a systematic attack taxonomy for LLM “app” platforms and apply it to OpenAI’s ChatGPT plugin framework, revealing concrete security vulnerabilities. These studies focus primarily on protocol description and security, leaving the scale and dynamics of MCP usage largely unexplored.
Agent Frameworks and Tool Libraries. Beyond platform-specific plugins, a variety of open-source frameworks provide standardized mechanisms for tool invocation by LLMs. LangChain and LlamaIndex define tool schemas for structured API calling (langchain2023, ; llamaindex2023, ). Hugging Face’s smolagents library (wolf2024smolagents, ) offers a lightweight Python API to build multi-step agentic workflows. Ge et al. conceptualize LLMs as an “AI Operating System” capable of hosting multiple agent applications (ge2023aios, ). These frameworks share MCP’s goal of formalizing LLM–tool interactions but their academic treatments remain largely descriptive, with little empirical evidence of real-world adoption.
@@ 10. Conclusion
This paper provides the first comprehensive measurement of the Model Context Protocol ecosystem. By collecting and analyzing 8,401 valid projects across six major markets, we shed light on its scale, security posture, and client interaction patterns. Our findings paint a mixed picture: while MCP has achieved rapid adoption, the ecosystem remains fragile, with over 50% of projects classified as low-value or abandoned. Servers show promising adoption of input validation frameworks in some languages, but also suffer from supply-chain monocultures, uneven maintenance, and risky exposure of sensitive APIs. Clients demonstrate convergence toward SSE as a dominant protocol, yet fragmentation persists, and most remain limited to single-server connections. These insights suggest that MCP is in a transitional stage—widely adopted in appearance but structurally fragile in practice. Moving forward, researchers and practitioners should explore methods for improving ecosystem sustainability, strengthening server security practices, and fostering interoperability among clients. Our dataset and framework aim to support these efforts, offering a foundation for future work on MCP standardization, governance, and security.

### v3

@@ 6.3. Measurement of Functionality and Implementation for Exposure Analysis
To comprehensively understand the MCP ecosystem from a security perspective, we measure servers along three functional dimensions: categories of exposed services, implementation languages, and API usage. Each of these dimensions provides a distinct lens on potential risks. First, server categories matter because certain functionalities such as authentication, personal data connectors, or proprietary model access, are inherently security-sensitive and expand the attack surface. Second, the implementation language is closely tied to security posture, since different language vary in their unsafe defaults (e.g., memory safety issues in C/C++ vs. stronger isolation in Rust). Third, sensitive API usage creates a broad attack surface for network, execution, and file-access threats.
@@
Figure 12. Sensitive API usage of MCP servers.
Table 5. Representative Sensitive APIs in MCP Server Environments.
API Type Threat Description Typical API Example
Network Request Allows making network connections, potentially enabling data exfiltration, SSRF attacks, or connections to malicious servers Python: requests.get, urllib.request.urlopen, socket.connect
Code Execution Permits execution of arbitrary code, allowing attackers to run malicious payloads within the application context Python: eval, exec
JavaScript: eval, Function()
System Command Executes system commands, possibly leading to unauthorized system control Python: os.system, subprocess.run, subprocess.Popen
JavaScript: child_process.exec, child_process.spawn
File Operation Enables reading or modifying files and directories, which could lead to data leakage or system compromise Python: open, os.remove, shutil.rmtree
JavaScript: fs.readFile, fs.writeFile
HTML Access Injects untrusted scripts or HTML into responses, potentially compromising client security JavaScript: document.write, element.innerHTML
To further assess the security posture of MCP servers, we conducted a static code scan to detect potential invocations of security‐sensitive APIs. The scan used a curated list of library functions commonly associated with high‐risk operations—such as network access, command execution, and dynamic code evaluation—as summarized in Table 5. Figure 12 summarizes the prevalence of five representative threat categories, with each bar showing the total number of servers containing at least one matched call and internal segments indicating their programming languages. In total, 4,095 servers were flagged as invoking at least one sensitive API. A few cases written in Rust or Ruby were also detected, but their counts are too small to be visible in Figure 12. The results show that Network Request APIs are by far the most prevalent, appearing in more than 2,700 servers across languages, dominated by JavaScript (2,160) and Python (512). Code Injection–related calls occur in 1,482 servers, again concentrated in JavaScript (1,260) with smaller but non‐negligible presence in Python (217). Command Execution primitives remain a major concern, with about 1,410 servers referencing OS‐level commands. Overall, these measurements indicate that MCP servers frequently include code patterns associated with security‐sensitive operations. While such matches do not necessarily imply active exploitation, they reveal a broad potential attack surface that warrants closer auditing. Regular dependency checks, sandboxed execution of network and file operations, and standardized authentication for sensitive API calls are essential to mitigate the risks surfaced by this analysis.
Answer to RQ2 (Security and Privacy Posture). Our measurements show that the MCP ecosystem is neither uniformly secure nor privacy-preserving, but instead exhibits a mix of encouraging practices and systemic risks. On the positive side, a large fraction of servers remain actively maintained (40.9% updated within the last 90 days), and many projects incorporate safeguards such as schema validation frameworks (e.g., pydantic, zod) that reduce the likelihood of injection or deserialization flaws. However, a substantial subset of servers include code that invokes security-sensitive APIs—for instance, network requests, command execution, or dynamic code evaluation—which, if misused or insufficiently sandboxed, could expose connected environments or external services to compromise. The language distribution amplifies this risk: more than 93% of MCP servers are implemented in JavaScript or Python, creating a supply-chain monoculture where vulnerabilities in popular npm or PyPI packages could cascade across thousands of deployments.
@@ 7. Measurement of MCP Clients
In this section, we conduct a systematic analysis of MCP clients, which function as the crucial interface between end-users, MCP servers, and the underlying LLMs that power model-driven tasks. By design, clients mediate the flow of requests and responses, orchestrating communication and shaping the overall user experience. Our measurement focuses on two complementary aspects. First, we examine communication protocols, which uncover whether client–server interactions are converging on a dominant mode or remain fragmented across competing designs. Protocol choice is not only a technical matter but also a signal of standardization, interoperability, and long-term ecosystem trajectory(toolfuzz). Second, we analyze cross-server usage, capturing whether clients are designed for single-purpose connections or instead support multi-server integration, which enables richer workflows, redundancy, and interoperability across heterogeneous backends. Together, these two perspectives highlight the balance between simplicity and extensibility in client design, and how this balance contributes to the ecosystem’s evolutionary path. To support this analysis, MCPCrawler collected a total of 341 valid MCP clients across four major markets. As part of a preprocessing step, we applied strict filtering criteria to remove 22 low-value entries, such as inactive forks, placeholder projects, or duplicates that added noise but no substantive functionality.
@@ 7.1. Measurement of Connection Protocols for Standardization Analysis
Figure 13. Transport mechanisms of MCP clients.
To characterize client connection patterns, we examined transport support among the 341 valid MCP clients, focusing on the three official communication mechanisms: stdio, Server-Sent Events (SSE), and streamable HTTP. Figure 13 presents the Venn diagram summarizing their overlap. The results reveal a conservative ecosystem with limited adoption of modern transports. Stdio remains the dominant mechanism, supported by nearly all clients (339 out of 341), reflecting a preference for local, synchronous communication inherited from early MCP prototypes. SSE, although officially deprecated since November 2024 according to the MCP specification(mcp2024transportupdate), continues to appear widely in co-implementations. In contrast, the officially recommended streamable HTTP transport shows limited adoption: as of August 2025, only 95 of 341 clients support it. This slow transition indicates that the client ecosystem has not kept pace with protocol evolution. Overall, MCP client development remains lagging behind specification updates. The persistence of legacy stdio and SSE communications, combined with the sluggish uptake of streamable HTTP, reflects substantial engineering inertia within the ecosystem, limiting interoperability and long-term maintainability.
@@ 7.2. Measurement of Client Connection Modes for Usage Dynamics
Answer to RQ3 (Client Connection Patterns in MCP Evolution). Our measurements reveal that MCP client connectivity remains fragmented despite protocol updates. Although SSE was formally deprecated in November 2024, it continues to dominate most implementations, while only 27.9% of clients have adopted the recommended streamable HTTP transport—evidence of slow migration and technical inertia. On the connection side, 80.9% of clients link to a single server, with 19.1% supporting multi-server integration, reflecting early but limited progress toward more complex and interoperable workflows. Overall, client design shows gradual movement toward standardization, yet persistent reliance on legacy mechanisms underscores the ecosystem’s uneven evolution.
@@ 8. Discussion
Ethical Considerations. Our study was conducted with explicit attention to research ethics. All data we analyzed were drawn from publicly accessible sources, including open repositories, market listings, and metadata exposed through official project listings. At no point did we attempt to bypass authentication mechanisms, crawl private datasets, or interact with servers in ways that could degrade their availability or stability. Importantly, downloading metadata and dependency information from Smithery or other MCP markets does not diminish the commercial or functional value of those platforms; rather, it mirrors the type of visibility already available to any developer or user browsing these resources. We did not attempt to probe or exploit server endpoints, ensuring that our methodology posed no threat to operational security. Beyond technical measures, we recognize broader ethical questions around ecosystem measurement, particularly the potential for exposing weaknesses that malicious actors could exploit. To mitigate this, our presentation emphasizes aggregate trends and structural observations, rather than detailed exploit paths.
@@ 9. Related Work
The MCP and the broader ecosystem of LLM plugin frameworks are emerging research frontiers that blend protocol design, marketplace dynamics, and security considerations. Because MCP itself was only recently introduced, existing literature remains sparse and fragmented. Previous large-scale measurements of software and application ecosystems (measurementgoogleplay; measurementminiapp) have demonstrated the effectiveness of empirical crawling and analysis in revealing structural patterns, maintenance issues, and dependency risks. Inspired by these approaches, we review two closely related lines of work: (i) early analyses of the MCP ecosystem, and (ii) open-source agent frameworks and tool libraries that share conceptual similarities with MCP but differ in deployment and measurement scope.
MCP Ecosystem Analyses. The Model Context Protocol (MCP) is a recent open JSON-RPC standard for connecting large language models (LLMs) to external tools and data sources (anthropic2024mcp). To date, only a handful of academic studies have investigated MCP. Li et al. (li2025urgentlyneedprivilegemanagement) conduct one of the first empirical analyses of real-world MCP-based plugins, revealing that network and system APIs dominate MCP servers and that low-download plugins frequently embed high-risk calls. Their work highlights the potential of MCP as a unifying protocol for LLM–tool integration but does not report deployment statistics or adoption rates. Similarly, Anthropic’s official technical documentation (anthropic2024mcp) defines the specification and server lifecycle but provides no usage analytics. Parallel to MCP, several works have analyzed OpenAI’s plugin-based ecosystems. Iqbal et al. (iqbal2024llmplatformsecurityapplying) propose a systematic attack taxonomy for LLM “app” platforms and apply it to OpenAI’s ChatGPT plugin framework, revealing concrete security vulnerabilities. Ehtesham et al. (surveyagentinteroperability) compare MCP with other agent interoperability protocols including ACP, A2A, and ANP, underscoring both its conceptual simplicity and its early-stage standardization challenges. Prior empirical work on large plugin ecosystems, such as WordPress, has shown that platform–plugin co-evolution strongly influences sustainability and innovation cycles (wordpress). Similar dynamics are beginning to emerge in MCP markets, where growth and fragmentation occur simultaneously. These studies focus primarily on protocol description and security, leaving the scale and dynamics of MCP usage largely unexplored.
Agent Frameworks and Tool Libraries. Beyond platform-specific plugins, a variety of open-source frameworks provide standardized mechanisms for tool invocation by LLMs. LangChain and LlamaIndex define tool schemas for structured API calling (langchain2023; llamaindex2023). Hugging Face’s smolagents library (wolf2024smolagents) offers a lightweight Python API to build multi-step agentic workflows. Ge et al. conceptualize LLMs as an “AI Operating System” capable of hosting multiple agent applications (ge2023aios). These frameworks share MCP’s goal of formalizing LLM–tool interactions but their academic treatments remain largely descriptive, with little empirical evidence of real-world adoption.
@@ 10. Conclusion
This paper provides the first comprehensive measurement of the Model Context Protocol ecosystem. By collecting and analyzing 8,401 valid projects across six major markets, we shed light on its scale, security posture, and client interaction patterns. Our findings paint a mixed picture: while MCP has achieved rapid adoption, the ecosystem remains fragile, with over 50% of projects classified as low-value or abandoned. Servers show promising adoption of input validation frameworks in some languages, but also suffer from supply-chain monocultures, uneven maintenance, and risky exposure of sensitive APIs. On the client side, SSE remains widely used even after its deprecation, and adoption of the newer streamable HTTP transport (27.9%) has been slow, reflecting the ecosystem’s gradual and uneven protocol evolution. These insights suggest that MCP is in a transitional stage—widely adopted in appearance but structurally fragile in practice. Moving forward, researchers and practitioners should explore methods for improving ecosystem sustainability, strengthening server security practices, and fostering interoperability among clients. Our dataset and framework aim to support these efforts, offering a foundation for future work on MCP standardization, governance, and security.